Based on historical data, the diameter of a ball bearing is normally distributed with a mean of 0.527 cm and a standard deviation of 0.009 cm. Suppose that a sample of 36 ball bearings are randomly selected. Determine the probability that the average diameter of a sampled ball bearing is greater than 0.530 cm.
a. 0.9772
b. 0.0228
c. 0.5062
d. 0.0559
Answer #1

Mean of sample (media%2F61a%2F61af8123-a3af-44e1-b6ca-55)= 0.527 cm

The standard deviation of the sample (media%2F5a1%2F5a163b15-0474-42c4-bce2-d4) = 0.009 cm

Number of observations per sample (n) = 36

Let Z be the value of the Z statistic from the standard normal distribution table.

Now let us consider an average diameter of 0.530cm.

0.530 = \mu +Z*\frac{\sigma }{\sqrt{n}}

0.009 0.530 = 0.527 + Z * V36

0.009 0.530-0.527 Z * 6

0.003 * б 0.009

Z = 2

The probability to the left side of the Z statistic(2) on a standard normal distribution table is 0.97725

The probability that the average diameter of a sampled ball bearing is greater than 0.530 cm = 1 - 0.97725

= 0.02275

Hence the correct answer is option B.